# Learning a Unified Degradation-aware Representation Model for Multi-modal Image Fusion

### v2

@@ Abstract
All-in-One Degradation-Aware Fusion Models (ADFMs), a class of multi-modal image fusion models, address complex scenes by mitigating degradations from source images and generating high-quality fused images. Mainstream ADFMs often rely on highly synthetic multi-modal multi-quality images for supervision, limiting their effectiveness in cross-modal and rare degradation scenarios. The inherent relationship among these multi-modal, multi-quality images of the same scene provides explicit supervision for training, but also raises above problems. To address these limitations, we present LURE, a Learning-driven Unified REpresentation model for infrared and visible Image Fusion, which is degradation-aware. LURE decouples multi-modal multi-quality data at the data level and recouples this relationship in a unified latent feature space (ULFS) by proposing a novel unified loss. This decoupling circumvents data-level limitations of prior models and allows leveraging real-world restoration datasets for training high-quality degradation-aware models, sidestepping above issues. To enhance text-image interaction, we refine image-text interaction and residual structures via Text-Guided Attention (TGA) and an inner residual structure. These enhances text’s spatial perception of images and preserve more visual details. Experiments show our method outperforms state-of-the-art (SOTA) methods across general fusion, degradation-aware fusion, and downstream tasks. The code will be publicly available.
Figure 1: Mechanisms of Existing Supervised Methods vs. Our Method. (a) Existing methods commonly utilize datasets constructed with quadruplets (i.e. 4 images, same scene, varying modalities & qualities), which introduces a series of issues. (b) Legend for symbols. (c) Our method disentangles this quadruplets at the data level and re-associates it in the feature space, thus achieving greater flexibility and versatility.
@@ 1 Introduction
Multi-modal image fusion, aims at creating high-quality fused images by integrating multi-modal sensor data; notably, infrared and visible image fusion is recognized as one of its crucial tasks [26, 16, 61]. It faces challenges due to degradations in visual information of different modalities. Detailed visible images degrade under varying weather and lighting, causing information loss. In contrast, infrared images, capturing thermal radiation in adverse environments, typically have lower resolution/contrast.[58, 56, 30]. Thus, effective degradation removal and complementary data integration are essential for accurate scene reconstruction [8].
A conventional two-step approach initially employs SOTA restoration models for modality-specific degradation removal, followed by fusion models for complementary information integration [56]. However, this strategy often results in domain shift due to the domain discrepancy between restoration and fusion models. Moreover, degradation combinations from both modalities, e.g., low light for visible images and low contrast for infrared images, will increase the challenge, rendering separate fusion model training for each degradation combination infeasible [25, 9, 62]..
Current methods address this problem in one step by training an All-in-One degradation-aware fusion model (ADFM). Mainstream ADFMs typically employ a supervised paradigm, relying on high-quality paired infrared and visible images as explicit supervisory signals to achieve high-quality fusion [56, 10]. However, these methods exhibit limited generalization to combined degradations and over-rely on synthetic data. Some unsupervised ADFMs address these issues [14]; for instance, DAFusion [48] utilizes contrastive learning and unpaired references, demonstrating improved generalization. Nevertheless, unsupervised methods often result in inferior fusion quality with detail loss due to the absence of explicit high-quality supervision.
To address these issues, analyzing image properties is crucial. Images are often characterized by modality and quality. Fig.1 (a) shows these dimensions form a coupled data quadruplet. Quality-wise, each modality requires paired low/high-quality images. Modality-wise, each image requires a cross-modality counterpart. This coupling causes two key issues: (1) High-quality restoration datasets become unusable due to lacking modality counterparts; (2) Sparse cross-modality degradation combinations lead to the “curse of dimensionality”. Furthermore, fusion datasets exhibit long-tailed degradation distributions, making crucial degradation synthesis challenging, e.g., fog needs depth maps [40, 1, 47].
The aforementioned problems stem from modality and quality coupling. While it facilitates end-to-end fusion and circumvents domain shift, it also introduces data-level challenges. As shown in Fig.1 (c), we addresses these by decoupling the quadruplets at the data-level, mapping data to a unified latent feature space (ULFS) that transforms needed data into binary tuples for restoration and fusion, thereby benefiting from a substantial amount of high-quality existing data, e.g., RESIDE [18], LOL [51], etc.
Fig.1 (c) shows these independent datasets sharing a latent space via mapping operation. Within this space, feature-level degradation combination becomes feasible, obviating data-level degradation combination concerns. However, the input-output residual connections of conventional restoration models, designed for better convergence, can also hinder latent representation learning by prioritizing degradation removal over source reconstruction [6, 57, 20, 7]. To counter this, we introduce a novel structure for efficient latent representation learning, named as inner-residual. Finally, we integrate text for flexible user control. To improve text-image feature interaction, we advance beyond simple channel weighting with Text-Guided Attention (TGA), enhancing text awareness spatially and channel-wise. The contributions of this paper can be summarized as follows:
@@
- An efficient inner residual structure and a text-image interaction mechanism are introduced to enhance the semantic interaction ability through textual conditioned latent feature representations.
@@ 2.1 Recent Image Fusion Methods
Mainstream image fusion models fall into three categories: generative models (e.g., FusionGAN, DDFM)[64, 33, 32], end-to-end approaches (e.g., U2Fusion)[27, 42, 63], and autoencoder-based methods (e.g., DenseFuse)[22, 24, 23]. Generative models fuse images via distribution approximation, yet face training instability and heuristic designs. End-to-end methods directly derive fused images, constrained by designed loss functions. Despite stable training, their generalization is limited by handcrafted loss functions. Early autoencoders lacked adaptability with hand-engineered fusion strategies[22], whereas current methods use learnable rules for better adaptation[23]. Autoencoder methods inherently decouple fusion rule and feature learning, facilitating versatile, degradation-aware paradigms.
@@ 2.2 All-in-One Degradation-aware Models
All-in-One Degradation-aware Models (ADM) address diverse degradations within a single model[10, 21, 20]. In image restoration field, ADMs benefit from abundant high-quality datasets and use techniques like prompt/continual learning[51, 18, 1].
However, in image fusion field, ADFMs face challenges due to long-tailed degradation distributions in fusion datasets, lacking sufficient data for robust degradation restoration[56, 31, 54, 41]. Unsupervised methods like contrastive learning exist but often yield lower quality due to lacking explicit supervision[48]. Current ADFMs still rely on existing fusion dataset degradations or synthetic ones, which may not match real degradation distributions[56, 48]. Some degradations require extra modality information for realistic synthesis (e.g., fog needs depth maps), which is unavailable in current datasets.
@@ 3.1 Problem Formulation
In this paper, low/high-quality images are denoted by sets $\mathcal{X}$/$\mathcal{Y}$. Infrared/visible images are denoted by sets $\mathcal{I}$/$\mathcal{V}$.
@@
For supervised method, samples are quadruplets $(x_{ir},y_{ir},x_{vi},y_{vi})$, where $ir$/$vi$ denote infrared/visible modalities, and $x$/$y$ are low/high-quality images. Define modality-paired sets: $\mathcal{I}_{g}$ (infrared), $\mathcal{V}_{g}$ (visible):
$\mathcal{I}_{g}=\{(x_{ir},y_{ir})\mid x_{ir}\in\mathcal{X},y_{ir}\in\mathcal{Y},y_{ir}=f(x_{ir})\}$ (1)
@@ 3.2 Data Decoupling and Feature Association
To address modality-quality coupling in data-level, a novel decoupling mechanism which still preserves inherent associations of source data is needed. We thus propose a ULFS ($\mathcal{Z}$) to re-establish these associations at feature level. Initially, we focus on image quality, disregarding modality, and define $\mathcal{Z}$ as follows:
@@
    Condition 1 ensures every degraded image maps to $\mathcal{Z}$.
@@
    Condition 2 ensures that for any latent feature $z\in\mathcal{Z}$, the posterior $P(c|z)$ distribution equals the prior $P(c)$ distribution for $c\in\mathcal{C}$.
@@
However, fusion rule learning still necessitates a modality-quality coupled dataset to associate first and second stage at feature level, incompletely decoupling dimensions. To achieve full decoupling, we introduce a pseudo-degradation task that auto-encodes the input image, thereby transferring this coupling to it.
@@ 3.2.1 Unified Latent Representation Learning
Figure 3: Schematic diagram of the two-stage training process. (a) In the first stage, a unified latent feature space is learned under the guidance of text. (b) In the second stage, all encoders are frozen, a fusion module is incorporated to learn a fusion strategy. ‘pseudo prompt’ refers to the dedicated prompt for the pseudo degradation and the detailed information of it can be found in supplementary materials.
@@
$\min_{\theta}\sum_{t=1}^{T}\sum_{t^{\prime}\neq t}^{T}\text{KL}(P^{(t)}_{\mathcal{Z}}(z)\parallel P^{(t^{\prime})}_{\mathcal{Z}}(z))$ (2)
[…]
To minimize Eq.2, Generative Adversarial Networks (GANs) with discriminators are commonly used for adversarial training [28, 37]. However, inherent distribution differences exist across image restoration datasets. Discriminators capture these discrepancies, leading to suboptimal representation learning for $f_{c}$ and sensitive GAN training, hindering model convergence.
To mitigate this issue, we introduce a pseudo-degradation task. This task, fundamentally image reconstruction, is not a genuine degradation. For this task, pseudo-degradation data pairs $(x,y)$ satisfy $x=y$, where $x\in\mathcal{X}$ and $y\in\mathcal{Y}$. We control this task via a description $c_{\mathrm{pd}}$ and align all degradation distributions to this pseudo-degradation task distribution. However, directly computing Kullback-Leibler divergence between tasks is challenging, requiring optimization objective transformation.
@@
$\mathcal{L_{\mathrm{unified}}}=\frac{1}{T}\sum_{t=1}^{T}\{1-\Gamma[f_{e}(X_{t},c_{t}),f_{e}(Y_{t},c_{\mathrm{pd}})]\}$ (3)
@@
Directly constraining feature relationships for $\mathcal{Z}$ learning is more stable than GAN-based adversarial methods. The pseudo-task serves as an image auto-encoding task to learn a fusion rule in the second stage and associate two stages at feature level. In practice, $f_{e}$ is a conditional image encoder.
@@ 3.3 Network Pipeline
In this section, the details of our model architecture are introduced, including Encoders ($E_{\mathrm{img}}$ and $E_{\mathrm{txt}}$), Fusion Module ($F$), and Decoder ($D$). Then, we describe two-stage training and loss function. Finally, the end-to-end inference for diverse scenarios is presented.
@@ 3.3.1 Encoder and Decoder Structure
As shown in Fig.3 (a), the encoder is bifurcated into image ($E_{\mathrm{img}}$) and text ($E_{\mathrm{txt}}$) encoders. $E_{\mathrm{txt}}$ employs Distilled BERT [13], with a Norm layer and a Linear layer to project its final output into a text feature vector. A classification head and cross-entropy loss ($\mathcal{L}_{\mathrm{text}}$) are used to align task categories for each task-specific text feature vector, an approach effective in InstructIR and superior to the CLIP text encoder [38].
$E_{\mathrm{img}}$ consists of multi-scale Encoder Layers. Learning $\mathcal{Z}$ requires effective mapping and reconstruction of the input image. Conventional image restoration models often utilize a residual structure ($\hat{y}=\mathcal{M}(x)+x$), where $\hat{y}$ represents the predicted high-quality image, $x$ is the low-quality input, and $\mathcal{M}$ is the restoration model [57, 20, 6]. This approach, however, prioritizes learning degraded regions over source image reconstruction, which may not suitable for learning $\mathcal{Z}$ in our framework.
Since directly eliminating the residual branch may lead to detail loss, we designed an inner residual structure to preserve details as much as possible. Each Encoder Layer (Fig.4) comprises four modules: BaseBlock, TGABlock (Text-Guided Attention Block), BottleNeck, and Linear. The TGABlock is combined with BaseBlock and BottleNeck to form two composite modules. Each module iterates $K^{i}_{tb}$ and $K^{i}_{tb}$ times in the $i$-th layer.
@@
Figure 4: Schematic diagrams of Encoder Layer and Text-Guided Attention (TGA). (a) Structure diagram of the Encoder Layer. (b) Structure diagram of TGA. “,$\odot$,$\oplus$,$\ominus$ , ⊕ , ⊖” represent Element-wise operations, and $\sigma$ represents the Sigmoid function.
@@
As shown in Fig.4 (a), at the $i$-th layer, TGABlock + BaseBlock extract modality-specific image information and initially integrate degradation information, guided by text. It’s output then branch into three paths: (1) direct downsampling to next layer; (2) degradation-specific information $\Phi_{d}^{i}$ extraction via TGABlock + BottleNeck; (3) task-relevant base information $\Phi_{t}^{i}$ filtering by a linear layer. Subsequently, a residual operation subtracts degradation-specific information $\Phi_{d}^{i}$ from task-relevant base information $\Phi_{t}^{i}$, yielding task-irrelevant latent representations: $z_{t}^{i}=\Phi_{t}^{i}-\Phi_{d}^{i}$. This inner residual connection provides a shorter path to decoder, preventing detailed information loss.
BaseBlock, consisting of NAFBlock [6] and Modality Embedding, serves to extract modality-specific basic information from images. The BottleNeck module, comprising a NAFBlock and a $1\times 1$ convolution, is designed to preserve text-relevant features and reduce redundant channels following the TGABlock.
For image-text interaction, the TGABlock utilizes a Transformer-like architecture, using GDFN (gated-Dconv FN) as the feed forward block [57]. In contrast to conventional methods that passively weight image channels with text [56, 36], which limits the global attention of text, our TGABlock can avoid this drawback.
As illustrated in Fig.4 (b), we use channel-weighted text information as the Query to proactively attend to global image features. These enhances text’s spatial perception of images and preserve more task-related visual information.
@@
Each decoder layer upsamples and simply employs few NAFBlocks to progressively reconstruct image details from $\mathcal{Z}$. For fusion task, a highly effective mechanism, cross attention, is used. Multiple Cross Attention modules are utilized at each scale.

### v3

# One Latent Space to Rule All Degradations: Unifying Restoration Knowledge for Image Fusion

@@ Abstract
All-in-One Degradation-Aware Fusion Models (ADFMs) as one of multi-modal image fusion models, which aims to address complex scenes by mitigating degradations from source images and generating high-quality fused images. Mainstream ADFMs rely on end-to-end learning and heavily synthesized datasets to achieve degradation awareness and fusion. This rough learning strategy and non-real world scenario dataset dependence often limit their upper-bound performance, leading to low-quality results. To address these limitations, we present LURE, a Learning-driven Unified REpresentation model for infrared and visible image fusion, which is degradation-aware. LURE learns a Unified Latent Feature Space (ULFS) to avoid the dependency on complex data formats inherent in previous end-to-end learning pipelines. It further improves image fusion quality by leveraging the intrinsic relationships between multi-modalities. A novel loss function is also proposed to drive the learning of unified latent representations more stable. More importantly, LURE seamlessly incorporates existing high-quality real-world image restoration datasets. To further enhance the model’s representation capability, we design a simple yet effective structure, termed internal residual block, to facilitate the learning of latent features. Experiments show our method outperforms state-of-the-art (SOTA) methods across general fusion, degradation-aware fusion, and downstream tasks. The code is available in the supplementary materials.
Figure 1: Illustration of the limitations in existing methods due to data-level issues. “Paired” denotes datasets containing corresponding infrared–visible image pairs.
@@ 1 Introduction
Multi-modal image fusion, aims at creating high-quality fused images by integrating multi-modal sensor data; notably, infrared and visible image fusion is recognized as one of its crucial tasks [bai2025task, karim2023current, zhang2025ddbfusion]. It faces challenges due to degradations in visual information of different modalities. Detailed visible images degrade under varying weather and lighting, causing information loss. In contrast, infrared images, capturing thermal radiation in adverse environments, typically have lower resolution/contrast.[zhang2021image, yi2024text, liang2022fusion]. Thus, effective degradation removal and complementary data integration are essential for accurate scene reconstruction [cheng2024fusionbooster].
Conventional two-step methods first restore and then fuse modalities [yi2024text], but suffer from domain shift between restoration and fusion. Moreover, diverse degradation combinations (e.g., visible low light and infrared low contrast) make separate fusion training infeasible [li2020mdlatlrr, cheng2025textfusion, zhao2023metafusion].
Current methods address this problem in one step by training an All-in-One degradation-aware fusion model (ADFM). Mainstream ADFMs typically employ a supervised paradigm, relying on high-quality paired infrared and visible images as explicit supervisory signals to achieve high-quality fusion [yi2024text, conde2024instructir]. This means that each training sample for these methods consists of a quadruplet of four images: two low-quality images of the same scene as input, and two corresponding high-quality images as supervision signals for end-to-end learning [yi2024text] [xu2020u2fusion] [jia2021llvip] [ha2017mfnet]. However, the datasets currently available to support the training of such fusion methods suffer from several critical deficiencies, which in turn constrain the fusion quality of existing models.
As illustrated in Fig.1, the infrared and visible modalities exhibit distinct types of degradation. These degradations are combined within the quadruplets required for training. Assuming there are $M$ types of degradation in the infrared modality and $N$ types in the visible modality, a total of $M\times N$ types of quadruplets would be necessary for comprehensive training. This leads to a combinatorial explosion and significantly increases the difficulty of dataset construction. Consequently, existing datasets heavily rely on synthetically generated data to simulate a wide array of degradation scenarios [xu2020u2fusion] [jia2021llvip] [ha2017mfnet].
For existing IVIF datasets (e.g., RoadScene [xu2020aaai], MSRS [Tang2022PIAFusion], M3FD [liu2022target]), degraded images are typically handled through two strategies. The forward construction method restores degraded images using SOTA restoration techniques or manual adjustment to obtain ground truth, while the backward construction method artificially adds degradations to clean images when real degraded samples are scarce. Forward construction is limited by the scarcity of real degraded scenes. In contrast, backward construction often introduces a distribution gap from real degradations—especially for rare cases. Current datasets follow a long-tail distribution, with most samples covering common degradations (e.g., low light, overexposure) and only a few representing rare but critical ones (e.g., dense fog).
To address these issues, we propose a representation learning method based on decoupling. First, in the initial stage, we leverage existing high-quality image restoration datasets for each modality independently. The objective is to learn a unified latent representation space that is agnostic to the specific type of degradation. In this space, the original degradation type cannot be inferred from an image’s representation, making it degradation-agnostic. It serves as a unified representation where all degraded images are mapped, effectively eliminating the influence of specific degradations. In the second stage, we build upon the features from this representation space to learn how to generate high-quality fused results—that is, to learn the fusion rules. Since all types of degradation are mapped to a common intermediate form, a single training process is sufficient to handle input images with any form of degradation. Importantly, for the first stage, we only require image restoration datasets, some of which provide high-quality supervision from real-world scenes (e.g., RESIDE [li2019benchmarking], LOL [wei2018deep]). For the second stage, we only need a standard image fusion dataset. This two-stage design enables us to circumvent the limitations of existing methods.
However, learning an effective image representation requires us to reconstruct the input image, ensuring that its representation preserves the critical information of the image. Existing image restoration models typically employ a skip connection [chen2022simple] [zamir2022restormer] [li2022all] [chen2023comparative]. from input to output to accelerate model convergence, allowing the model to focus on the degraded regions of the image. While this architecture hinders the learning of a comprehensive image representation, removing it leads to slower convergence. Therefore, we adapt this residual learning mechanism by shifting from the conventional pixel-level residual to a feature-level residual, proposing an “inner residual structure” to resolve the aforementioned issue. The contributions of this paper can be summarized as follows:
@@
- An efficient inner residual structure is introduced to enhance the model’s information retention while ensuring its ability to learn complete latent features.
@@ 2.1 Recent Image Fusion Methods
Mainstream image fusion models fall into three categories: generative models (e.g., FusionGAN, DDFM)[zhao2023ddfm, ma2020ganmcc, ma2019fusiongan], end-to-end approaches (e.g., U2Fusion)[li2025conti, wang2025highlight, zhao2023cddfuse], and autoencoder-based methods (e.g., DenseFuse)[li2018densefuse, li2020nestfuse, li2024crossfuse]. Generative models fuse images via distribution approximation, yet face training instability and heuristic designs. End-to-end methods directly derive fused images, constrained by designed loss functions.
Despite stable training, their generalization is limited by handcrafted loss functions. Early autoencoders lacked adaptability with hand-engineered fusion strategies[li2018densefuse], whereas current methods use learnable rules for better adaptation [li2024crossfuse]. Autoencoder methods inherently decouple fusion rule and feature learning, facilitating versatile, degradation-aware paradigms.
@@ 2.2 All-in-One Degradation-aware Models
All-in-One Degradation-aware Models (ADM) address diverse degradations within a single model[conde2024instructir, li2024promptcir, li2022all]. In image restoration field, ADMs benefit from abundant high-quality datasets and use techniques like prompt/continual learning[wei2018deep, li2019benchmarking, afifi2021learning].
However, in image fusion field, ADFMs face challenges due to long-tailed degradation distributions in fusion datasets, lacking sufficient data for robust degradation restoration[yi2024text, liu2022target, xu2020aaai, Tang2022PIAFusion]. Unsupervised methods like contrastive learning exist but often yield lower quality due to lacking explicit supervision[wang2025degradation]. Current ADFMs still rely on existing fusion dataset degradations or synthetic ones, which may not match real degradation distributions[yi2024text, wang2025degradation]. Some degradations require extra modality information for realistic synthesis (e.g., fog needs depth maps), which is unavailable in current datasets.
@@ 3.1 Problem Formulation
In this paper, low/high-quality images are denoted by sets $\mathcal{X}$$\mathcal{Y}$. Infrared/visible images are denoted by sets $\mathcal{I}$$\mathcal{V}$.
@@
For supervised method, samples are quadruplets $(x_{ir},y_{ir},x_{vi},y_{vi})$, where irir/vivi denote infrared/visible modalities, and xx/$y$ are low/high-quality images. Define modality-paired sets: $\mathcal{I}_{g}$ (infrared), $\mathcal{V}_{g}$ (visible):
$\mathcal{I}_{g}=\{(x_{ir},y_{ir})\mid x_{ir}\in\mathcal{X},y_{ir}\in\mathcal{Y},y_{ir}=f(x_{ir})\},$ (1)
@@ 3.2 Data Decoupling and Feature Association
To address modality-quality coupling in data-level, a novel decoupling mechanism which still preserves inherent associations of source data is needed. We thus propose a ULFS ($\mathcal{Z}$) to re-establish these associations at feature level. We define $\mathcal{Z}$ as follows:
@@
    Condition I ensures every degraded image maps to $\mathcal{Z}$.
@@
    Condition II ensures that for any latent feature $z\in\mathcal{Z}$, the posterior $P(c|z)$ distribution equals the prior $P(c)$ distribution for $c\in\mathcal{C}$.
@@ 3.2.1 Unified Latent Representation Learning
Figure 3: Two-stage training overview. (a) Stage I learns a unified latent space guided by text. (b) Stage II freezes encoders and trains a fusion module for strategy learning. “Pseudo prompt” denotes the prompt for pseudo degradation (details in Supplementary).
@@
$\min_{\theta}\sum_{t=1}^{T}\sum_{t^{\prime}\neq t}^{T}\text{KL}(P^{(t)}_{\mathcal{Z}}(z)\parallel P^{(t^{\prime})}_{\mathcal{Z}}(z)),$ (2)
where, $\text{KL}(\cdot\parallel\cdot)$ denotes the Kullback-Leibler divergence, and $P^{(t)}_{\mathcal{Z}}(z)$ represents the distribution of the $t$-th degradation in $\mathcal{Z}$.
To minimize Eq.2, Generative Adversarial Networks (GANs) with discriminators are commonly used for adversarial training [li2023learning, poirier2023robust]. However, inherent distribution differences exist across image restoration datasets. As the discriminators tend to capture these discrepancies, the GAN training process becomes unstable and struggles to converge.
To mitigate this issue, we introduce a pseudo-degradation task. This task, fundamentally image reconstruction, is not a genuine degradation. For this task, pseudo-degradation data pairs $(x,y)$ satisfy $x=y$, where $x\in\mathcal{X}$ and $y\in\mathcal{Y}$. We control this task via a description $c_{\mathrm{pd}}$ and align all degradation distributions to this pseudo-degradation task distribution.
@@
$\mathcal{L_{\mathrm{unified}}}=\frac{1}{T}\sum_{t=1}^{T}\{1-\Gamma[f_{e}(X_{t},c_{t}),f_{e}(Y_{t},c_{\mathrm{pd}})]\},$ (3)
@@
It is worth emphasizing that, in practice, the distributional discrepancies among different datasets still prevent Condition II in the ULFS definition from being perfectly satisfied. Consequently, the learned latent space can only approximately meet this condition. Nevertheless, directly constraining feature relationships for $\mathcal{Z}$ learning proves to be more stable than GAN-based adversarial approaches. The pseudo-task acts as an image auto-encoding process to learn the fusion rule in the second stage and to bridge the two stages at the feature level. In practice, $f_{e}$ functions as a conditional image encoder.
@@ 3.3.1 Encoder and Decoder Structure
As shown in Fig.3 (a), the encoder is bifurcated into image ($E_{\mathrm{img}}$) and text ($E_{\mathrm{txt}}$) encoders. $E_{\mathrm{txt}}$ employs Distilled BERT [devlin2019bert], with a Norm layer and a Linear layer to project its final output into a text feature vector. A classification head and cross-entropy loss ($\mathcal{L}_{\mathrm{text}}$) are used to align task categories for each task-specific text feature vector, an approach effective in InstructIR [oh2024instructir] and superior to the CLIP text encoder [radford2021learning].
$E_{\mathrm{img}}$ comprises multi-scale Encoder Layers for mapping and reconstructing input images to learn $\mathcal{Z}$. Unlike conventional restoration models using residual structures ($\hat{y}=\mathcal{M}(x)+x$) [zamir2022restormer, li2022all, chen2022simple], which emphasize degraded-region correction over source reconstruction, our framework requires balanced feature learning. To avoid detail loss from removing the residual path, we introduce an inner residual structure to preserve fine information. Each Encoder Layer (Fig.4) includes four modules—BaseBlock, TGABlock (Text-Guided Attention Block), BottleNeck, and Linear—where TGABlock is integrated with BaseBlock and BottleNeck to form two composite units, iterated $K^{i}_{tb}$ and $K^{i}_{tb}$ times in the $i$-th layer.
@@
Figure 4: Schematic diagrams of Encoder Layer and Text-Guided Attention (TGA). (a) Structure diagram of the Encoder Layer. (b) Structure diagram of TGA. $\odot,\oplus,\ominus$” represent Element-wise operations, and $\sigma$ represents the Sigmoid function.
@@
As shown in Fig.4, at the $i$-th layer, TGABlock + BaseBlock extract modality-specific features and integrate degradation information guided by text. The output branches into three paths: (1) downsampling to the next layer, (2) TGABlock + BottleNeck for degradation-specific feature extraction $\Phi_{d}^{i}$, and (3) a linear layer for task-relevant base information $\Phi_{t}^{i}$. A residual subtraction yields task-irrelevant latent representations $z_{t}^{i}=\Phi_{t}^{i}-\Phi_{d}^{i}$, providing a shorter decoding path and mitigating detail loss.
BaseBlock (NAFBlock [chen2022simple] + Modality Embedding) extracts modality-aware basic features, while BottleNeck (NAFBlock + $1\times 1$ conv) retains text-relevant cues and reduces redundancy. The TGABlock adopts a Transformer-like design with GDFN [zamir2022restormer] as its feed-forward block. Unlike passive channel-weighting schemes [yi2024text, oh2024instructir], TGABlock employs channel-weighted text features as Queries to actively attend to global image features, enhancing text-guided spatial awareness and preserving task-related details.
@@
Each decoder layer then upsamples and refines details from $\mathcal{Z}$ using a few NAFBlocks, while the fusion stage incorporates multi-scale Cross Attention modules for effective multimodal interaction.